The side scale factor of 2 rectangles is 3:5.
Find the Area of rectangle ABCD, if smaller rectangle EFGH area is 32

Answer by CubeyThePenguin(3113)

side scale factor: 3/5
area factor: (3/5)^2 = 9/25

9/25 = 32/x

x = (25 * 32)/9 = 88 8/9